\begin{thebibliography}{54}
\expandafter\ifx\csname natexlab\endcsname\relax\def\natexlab#1{#1}\fi

\bibitem[{Caswell et~al.(2019)Caswell, Chelba, and
  Grangier}]{caswell-etal-2019-tagged}
Isaac Caswell, Ciprian Chelba, and David Grangier. 2019.
\newblock Tagged back-translation.
\newblock In \emph{WMT}.

\bibitem[{Chen et~al.(2015)Chen, Zhang, Li, and Wang}]{chen2015bilingual}
Chao-Ying Chen, John~Xuexin Zhang, Li~Li, and Ruiming Wang. 2015.
\newblock Bilingual memory representations in less fluent chinese--english
  bilinguals: An event-related potential study.
\newblock \emph{Psychological reports}.

\bibitem[{Cohn et~al.(2016)Cohn, Hoang, Vymolova, Yao, Dyer, and
  Haffari}]{cohn2016incorporating}
Trevor Cohn, Cong Duy~Vu Hoang, Ekaterina Vymolova, Kaisheng Yao, Chris Dyer,
  and Gholamreza Haffari. 2016.
\newblock Incorporating structural alignment biases into an attentional neural
  translation model.
\newblock In \emph{NAACL}.

\bibitem[{Collins et~al.(2005)Collins, Koehn, and
  Ku{\v{c}}erov{\'a}}]{collins2005clause}
Michael Collins, Philipp Koehn, and Ivona Ku{\v{c}}erov{\'a}. 2005.
\newblock Clause restructuring for statistical machine translation.
\newblock In \emph{ACL}.

\bibitem[{Conneau and Lample(2019)}]{lample2019cross}
Alexis Conneau and Guillaume Lample. 2019.
\newblock Cross-lingual language model pretraining.
\newblock In \emph{NeurIPS}.

\bibitem[{Ding and Tao(2019)}]{ding2019wmt}
Liang Ding and Dacheng Tao. 2019.
\newblock The university of sydney’s machine translation system for wmt19.
\newblock In \emph{WMT}.

\bibitem[{Ding et~al.(2021{\natexlab{a}})Ding, Wang, Liu, Wong, Tao, and
  Tu}]{Ding2021acl2}
Liang Ding, Longyue Wang, Xuebo Liu, Derek~F. Wong, Dacheng Tao, and Zhaopeng
  Tu. 2021{\natexlab{a}}.
\newblock Progressive multi-granularity training for non-autoregressive
  translation.
\newblock In \emph{findings of ACL}.

\bibitem[{Ding et~al.(2021{\natexlab{b}})Ding, Wang, Liu, Wong, Tao, and
  Tu}]{Ding2021acl1}
Liang Ding, Longyue Wang, Xuebo Liu, Derek~F. Wong, Dacheng Tao, and Zhaopeng
  Tu. 2021{\natexlab{b}}.
\newblock Rejuvenating low-frequency words: Making the most of parallel data in
  non-autoregressive translation.
\newblock In \emph{ACL}.

\bibitem[{Ding et~al.(2021{\natexlab{c}})Ding, Wang, Liu, Wong, Tao, and
  Tu}]{ding2021iclr}
Liang Ding, Longyue Wang, Xuebo Liu, Derek~F. Wong, Dacheng Tao, and Zhaopeng
  Tu. 2021{\natexlab{c}}.
\newblock Understanding and improving lexical choice in non-autoregressive
  translation.
\newblock In \emph{ICLR}.

\bibitem[{Ding et~al.(2020{\natexlab{a}})Ding, Wang, and Tao}]{ding2020acl}
Liang Ding, Longyue Wang, and Dacheng Tao. 2020{\natexlab{a}}.
\newblock Self-attention with cross-lingual position representation.
\newblock In \emph{ACL}.

\bibitem[{Ding et~al.(2020{\natexlab{b}})Ding, Wang, Wu, Tao, and
  Tu}]{Ding2020coling}
Liang Ding, Longyue Wang, Di~Wu, Dacheng Tao, and Zhaopeng Tu.
  2020{\natexlab{b}}.
\newblock Context-aware cross-attention for non-autoregressive translation.
\newblock In \emph{COLING}.

\bibitem[{Ding et~al.(2021{\natexlab{d}})Ding, Wu, and Tao}]{Ding2021iwslt}
Liang Ding, Di~Wu, and Dacheng Tao. 2021{\natexlab{d}}.
\newblock The usyd-jd speech translation system for iwslt2021.
\newblock In \emph{IWSLT}.

\bibitem[{Dworin(2003)}]{dworin2003insights}
Joel~E Dworin. 2003.
\newblock Insights into biliteracy development: Toward a bidirectional theory
  of bilingual pedagogy.
\newblock \emph{Journal of Hispanic Higher Education}, 2(2):171--186.

\bibitem[{Edunov et~al.(2018)Edunov, Ott, Auli, and
  Grangier}]{edunov2018understanding}
Sergey Edunov, Myle Ott, Michael Auli, and David Grangier. 2018.
\newblock Understanding back-translation at scale.
\newblock In \emph{EMNLP}.

\bibitem[{Garg et~al.(2019)Garg, Peitz, Nallasamy, and
  Paulik}]{garg2019jointly}
Sarthak Garg, Stephan Peitz, Udhyakumar Nallasamy, and Matthias Paulik. 2019.
\newblock Jointly learning to align and translate with transformer models.
\newblock In \emph{EMNLP}.

\bibitem[{Gehring et~al.(2017)Gehring, Auli, Grangier, Yarats, and
  Dauphin}]{gehring2017convolutional}
Jonas Gehring, Michael Auli, David Grangier, Denis Yarats, and Yann~N Dauphin.
  2017.
\newblock Convolutional sequence to sequence learning.
\newblock In \emph{ICML}.

\bibitem[{Gu et~al.(2018)Gu, Bradbury, Xiong, Li, and Socher}]{gu2018non}
Jiatao Gu, James Bradbury, Caiming Xiong, Victor~OK Li, and Richard Socher.
  2018.
\newblock Non-autoregressive neural machine translation.
\newblock In \emph{ICLR}.

\bibitem[{Ha et~al.(2016)Ha, Niehues, and Waibel}]{hatoward}
Thanh-Le Ha, Jan Niehues, and Alex Waibel. 2016.
\newblock Toward multilingual neural machine translation with universal encoder
  and decoder.
\newblock In \emph{IWSLT}.

\bibitem[{Hassan et~al.(2018)Hassan, Aue, Chen, Chowdhary, Clark, Federmann,
  Huang, Junczys-Dowmunt, Lewis, Li et~al.}]{hassan2018achieving}
Hany Hassan, Anthony Aue, Chang Chen, Vishal Chowdhary, Jonathan Clark,
  Christian Federmann, Xuedong Huang, Marcin Junczys-Dowmunt, William Lewis,
  Mu~Li, et~al. 2018.
\newblock Achieving human parity on automatic chinese to english news
  translation.
\newblock In \emph{arXiv}.

\bibitem[{He et~al.(2016)He, Xia, Qin, Wang, Yu, Liu, and Ma}]{he2016dual}
Di~He, Yingce Xia, Tao Qin, Liwei Wang, Nenghai Yu, Tie-Yan Liu, and Wei-Ying
  Ma. 2016.
\newblock Dual learning for machine translation.
\newblock \emph{NeurIPS}.

\bibitem[{He et~al.(2018)He, Tan, Xia, He, Qin, Chen, and Liu}]{he2018layer}
Tianyu He, Xu~Tan, Yingce Xia, Di~He, Tao Qin, Zhibo Chen, and Tie-Yan Liu.
  2018.
\newblock Layer-wise coordination between encoder and decoder for neural
  machine translation.
\newblock In \emph{NeurIPS}.

\bibitem[{Hendrycks et~al.(2019)Hendrycks, Lee, and
  Mazeika}]{hendrycks2019using}
Dan Hendrycks, Kimin Lee, and Mantas Mazeika. 2019.
\newblock Using pre-training can improve model robustness and uncertainty.
\newblock In \emph{ICML}.

\bibitem[{Kim and Rush(2016)}]{kim-rush-2016-sequence}
Yoon Kim and Alexander~M. Rush. 2016.
\newblock Sequence-level knowledge distillation.
\newblock In \emph{EMNLP}.

\bibitem[{Kingma and Ba(2015)}]{kingma2015adam}
Diederik~P. Kingma and Jimmy Ba. 2015.
\newblock Adam: A method for stochastic optimization.
\newblock In \emph{ICLR}.

\bibitem[{Koehn and Knowles(2017)}]{koehn2017six}
Philipp Koehn and Rebecca Knowles. 2017.
\newblock Six challenges for neural machine translation.
\newblock In \emph{WNMT}.

\bibitem[{Lample et~al.(2018)Lample, Ott, Conneau, Denoyer, and
  Ranzato}]{lample2018phrase}
Guillaume Lample, Myle Ott, Alexis Conneau, Ludovic Denoyer, and Marc’Aurelio
  Ranzato. 2018.
\newblock Phrase-based \& neural unsupervised machine translation.
\newblock In \emph{EMNLP}.

\bibitem[{Levinboim et~al.(2015)Levinboim, Vaswani, and
  Chiang}]{levinboim-etal-2015-model}
Tomer Levinboim, Ashish Vaswani, and David Chiang. 2015.
\newblock Model invertibility regularization: Sequence alignment with or
  without parallel data.
\newblock In \emph{NAACL}.

\bibitem[{Li et~al.(2019)Li, Li, Xu, Lin, Liu, Liu, Wang, Zhang, Xu, Wang
  et~al.}]{li2019niutrans}
Bei Li, Yinqiao Li, Chen Xu, Ye~Lin, Jiqiang Liu, Hui Liu, Ziyang Wang, Yuhao
  Zhang, Nuo Xu, Zeyang Wang, et~al. 2019.
\newblock The niutrans machine translation systems for wmt19.
\newblock In \emph{WMT}.

\bibitem[{Liang et~al.(2007)Liang, Klein, and
  Jordan}]{Liang2007AgreementBasedL}
P.~Liang, D.~Klein, and Michael~I. Jordan. 2007.
\newblock Agreement-based learning.
\newblock In \emph{NeurIPS}.

\bibitem[{Lin et~al.(2020)Lin, Pan, Wang, Qiu, Feng, Zhou, and Li}]{lin2020pre}
Zehui Lin, Xiao Pan, Mingxuan Wang, Xipeng Qiu, Jiangtao Feng, Hao Zhou, and
  Lei Li. 2020.
\newblock Pre-training multilingual neural machine translation by leveraging
  alignment information.
\newblock In \emph{EMNLP}.

\bibitem[{Liu et~al.(2020{\natexlab{a}})Liu, Lai, Wong, and Chao}]{liu2020acl}
Xuebo Liu, Houtim Lai, Derek~F. Wong, and Lidia~S. Chao. 2020{\natexlab{a}}.
\newblock Norm-based curriculum learning for neural machine translation.
\newblock In \emph{ACL}.

\bibitem[{Liu et~al.(2021)Liu, Wang, Wong, Ding, Chao, and Tu}]{liu2021emnlp}
Xuebo Liu, Longyue Wang, Derek~F. Wong, Liang Ding, Lidia~S. Chao, and Zhaopeng
  Tu. 2021.
\newblock On the complementarity between pre-training and back-translation for
  neural machine translation.
\newblock In \emph{findings of EMNLP}.

\bibitem[{Liu et~al.(2020{\natexlab{b}})Liu, Gu, Goyal, Li, Edunov,
  Ghazvininejad, Lewis, and Zettlemoyer}]{Liu2020mbart}
Yinhan Liu, Jiatao Gu, Naman Goyal, Xian Li, Sergey Edunov, Marjan
  Ghazvininejad, Mike Lewis, and Luke Zettlemoyer. 2020{\natexlab{b}}.
\newblock Multilingual denoising pre-training for neural machine translation.
\newblock \emph{Transactions of the Association for Computational Linguistics}.

\bibitem[{Luong et~al.(2015)Luong, Pham, and Manning}]{luong2015effective}
Minh-Thang Luong, Hieu Pham, and Christopher~D Manning. 2015.
\newblock Effective approaches to attention-based neural machine translation.
\newblock In \emph{EMNLP}.

\bibitem[{Mathis et~al.(2021)Mathis, Biasi, Schneider, Yuksekgonul, Rogers,
  Bethge, and Mathis}]{mathis2021pretraining}
Alexander Mathis, Thomas Biasi, Steffen Schneider, Mert Yuksekgonul, Byron
  Rogers, Matthias Bethge, and Mackenzie~W Mathis. 2021.
\newblock Pretraining boosts out-of-domain robustness for pose estimation.
\newblock In \emph{WACV}.

\bibitem[{Morishita et~al.(2017)Morishita, Suzuki, and
  Nagata}]{morishita2017ntt}
Makoto Morishita, Jun Suzuki, and Masaaki Nagata. 2017.
\newblock Ntt neural machine translation systems at wat 2017.
\newblock In \emph{IJCNLP}.

\bibitem[{Nguyen et~al.(2020)Nguyen, Joty, Kui, and Aw}]{nguyen2019data}
Xuan-Phi Nguyen, Shafiq Joty, Wu~Kui, and Ai~Ti Aw. 2020.
\newblock Data diversification: A simple strategy for neural machine
  translation.
\newblock In \emph{NeurIPS}.

\bibitem[{Och and Ney(2003)}]{och2003systematic}
Franz~Josef Och and Hermann Ney. 2003.
\newblock A systematic comparison of various statistical alignment models.
\newblock \emph{Computational linguistics}.

\bibitem[{Papineni et~al.(2002)Papineni, Roukos, Ward, and
  Zhu}]{papineni2002bleu}
Kishore Papineni, Salim Roukos, Todd Ward, and Wei-Jing Zhu. 2002.
\newblock Bleu: a method for automatic evaluation of machine translation.
\newblock In \emph{ACL}.

\bibitem[{Pavlenko and Jarvis(2002)}]{pavlenko2002bidirectional}
Aneta Pavlenko and Scott Jarvis. 2002.
\newblock Bidirectional transfer.
\newblock \emph{Applied linguistics}.

\bibitem[{Post(2018)}]{post-2018-call}
Matt Post. 2018.
\newblock A call for clarity in reporting {BLEU} scores.
\newblock In \emph{WMT}.

\bibitem[{Sennrich et~al.(2016{\natexlab{a}})Sennrich, Haddow, and
  Birch}]{sennrich-etal-2016-improving}
Rico Sennrich, Barry Haddow, and Alexandra Birch. 2016{\natexlab{a}}.
\newblock Improving neural machine translation models with monolingual data.
\newblock In \emph{ACL}.

\bibitem[{Sennrich et~al.(2016{\natexlab{b}})Sennrich, Haddow, and
  Birch}]{Sennrich:BPE}
Rico Sennrich, Barry Haddow, and Alexandra Birch. 2016{\natexlab{b}}.
\newblock Neural machine translation of rare words with subword units.
\newblock In \emph{ACL}.

\bibitem[{Sennrich and Zhang(2019)}]{sennrich2019revisiting}
Rico Sennrich and Biao Zhang. 2019.
\newblock Revisiting low-resource neural machine translation: A case study.
\newblock In \emph{ACL}.

\bibitem[{Tang et~al.(2019)Tang, Sennrich, and Nivre}]{tang2019understanding}
Gongbo Tang, Rico Sennrich, and Joakim Nivre. 2019.
\newblock Understanding neural machine translation by simplification: The case
  of encoder-free models.
\newblock In \emph{RANLP}.

\bibitem[{Vaswani et~al.(2017)Vaswani, Shazeer, Parmar, Uszkoreit, Jones,
  Gomez, Kaiser, and Polosukhin}]{transformer}
Ashish Vaswani, Noam Shazeer, Niki Parmar, Jakob Uszkoreit, Llion Jones,
  Aidan~N Gomez, Lukasz Kaiser, and Illia Polosukhin. 2017.
\newblock Attention is all you need.
\newblock In \emph{NeurIPS}.

\bibitem[{Wang et~al.(2020)Wang, Tu, Wang, Ding, Ding, and
  Shi}]{Wang2020TencentAL}
Longyue Wang, Zhaopeng Tu, Xing Wang, Li~Ding, Liang Ding, and Shuming Shi.
  2020.
\newblock Tencent ai lab machine translation systems for wmt20 chat translation
  task.
\newblock In \emph{WMT}.

\bibitem[{Wu et~al.(2019{\natexlab{a}})Wu, Fan, Baevski, Dauphin, and
  Auli}]{wu2019pay}
Felix Wu, Angela Fan, Alexei Baevski, Yann~N Dauphin, and Michael Auli.
  2019{\natexlab{a}}.
\newblock Pay less attention with lightweight and dynamic convolutions.
\newblock In \emph{ICLR}.

\bibitem[{Wu et~al.(2019{\natexlab{b}})Wu, Wang, Xia, Tao, Lai, and
  Liu}]{wu2019exploiting}
Lijun Wu, Yiren Wang, Yingce Xia, QIN Tao, Jianhuang Lai, and Tie-Yan Liu.
  2019{\natexlab{b}}.
\newblock Exploiting monolingual data at scale for neural machine translation.
\newblock In \emph{EMNLP}.

\bibitem[{Wu et~al.(2016)Wu, Schuster, Chen, Le, Norouzi, Macherey, Krikun,
  Cao, Gao, Macherey et~al.}]{wu2016google}
Yonghui Wu, Mike Schuster, Zhifeng Chen, Quoc~V Le, Mohammad Norouzi, Wolfgang
  Macherey, Maxim Krikun, Yuan Cao, Qin Gao, Klaus Macherey, et~al. 2016.
\newblock Google's neural machine translation system: Bridging the gap between
  human and machine translation.
\newblock \emph{arXiv}.

\bibitem[{Yang et~al.(2020)Yang, Hu, Han, Huang, and Ju}]{yang2020csp}
Zhen Yang, Bojie Hu, Ambyera Han, Shen Huang, and Qi~Ju. 2020.
\newblock Csp: Code-switching pre-training for neural machine translation.
\newblock In \emph{EMNLP}.

\bibitem[{Zhan et~al.(2021)Zhan, Liu, Wong, and Chao}]{zhan2021meta}
Runzhe Zhan, Xuebo Liu, Derek~F Wong, and Lidia~S Chao. 2021.
\newblock Meta-curriculum learning for domain adaptation in neural machine
  translation.
\newblock In \emph{AAAI}.

\bibitem[{Zheng et~al.(2019)Zheng, Zhou, Huang, Li, Dai, and
  Chen}]{zheng2019mirror}
Zaixiang Zheng, Hao Zhou, Shujian Huang, Lei Li, Xin-Yu Dai, and Jiajun Chen.
  2019.
\newblock Mirror-generative neural machine translation.
\newblock In \emph{ICLR}.

\bibitem[{Zhou et~al.(2021)Zhou, Ding, Duh, Sasano, and Takeda}]{zhou2021self}
Lei Zhou, Liang Ding, Kevin Duh, Ryohei Sasano, and Koichi Takeda. 2021.
\newblock Self-guided curriculum learning for neural machine translation.
\newblock In \emph{IWSLT}.

\end{thebibliography}

\pdfoutput=1

\documentclass[11pt]{article}

\usepackage[]{emnlp2021}
\usepackage{enumitem}
\usepackage{times}
\usepackage{latexsym}
\usepackage{xspace}
\usepackage[T1]{fontenc}
\usepackage{graphicx}
\usepackage[utf8]{inputenc}

\usepackage{microtype}
\usepackage[utf8]{inputenc} 
\usepackage[T1]{fontenc}    
\usepackage{hyperref}       
\usepackage{url}            
\usepackage{booktabs,colortbl}       
\usepackage{amsfonts}       
\usepackage{nicefrac}       
\usepackage{microtype}      
\usepackage{xcolor}         
\usepackage{CJKutf8}
\usepackage{balance}
\usepackage{bm}
\usepackage{amssymb}
\usepackage{url}
\usepackage{makecell}
\usepackage{colortbl}
\usepackage{textcomp}
\usepackage{CJK}
\usepackage{enumitem}
\usepackage{amsfonts}
\usepackage{caption}
\usepackage{float}
\usepackage{multirow}
\usepackage{color}
\usepackage{arydshln}
\usepackage{wrapfig}
\usepackage{graphicx}
\usepackage{amsmath}
\usepackage{enumitem}
\usepackage{graphics}
\usepackage{graphicx}
\usepackage{subfig}

\usepackage{url}            
\usepackage{booktabs}       
\usepackage{amsfonts}       
\usepackage{nicefrac}       
\usepackage{microtype}      
\usepackage{amsmath}
\usepackage{makecell}
\usepackage{caption}
\usepackage{float}
\usepackage{multirow}
\usepackage{color}
\usepackage{arydshln}
\definecolor{liam}{RGB}{220,220,220}
\DeclareMathOperator*{\argmax}{arg\,max}
\renewcommand\vec[1]{\overrightarrow{#1}}
\newcommand\cev[1]{\overleftarrow{#1}}
\newcommand\bid[1]{\overleftrightarrow{#1}}

\title{Improving Neural Machine Translation by Bidirectional Training}

\author{
Liang Ding\\
The University of Sydney\\
\normalsize \texttt{ldin3097@sydney.edu.au}\\
\And
Di Wu
\\
Peking University\\
\normalsize \texttt{inbath@163.com}\\
\And
Dacheng Tao\\
JD Explore Academy, JD.com\\
\normalsize \texttt{dacheng.tao@gmail.com}}

\begin{document}
\maketitle
\begin{abstract}
We present a simple and effective pretraining strategy -- bidirectional training (BiT) for neural machine translation. Specifically, we bidirectionally update the model parameters at the early stage and then tune the model normally. To achieve bidirectional updating, we simply reconstruct the training samples from ``src$\rightarrow$tgt'' to  ``src+tgt$\rightarrow$tgt+src'' without any complicated model modifications.  Notably, our approach does not increase any parameters or training steps, requiring the parallel data merely. Experimental results show that BiT pushes the SOTA neural machine translation performance across 15 translation tasks on 8 language pairs (data sizes range from 160K to 38M) significantly higher. Encouragingly, our proposed model can complement existing data manipulation strategies, i.e. back translation, data distillation and data diversification. Extensive analyses show that our approach functions as a novel bilingual code-switcher, obtaining better bilingual alignment.
\end{abstract}

\section{Introduction}
Recent years have seen a surge of interest in neural machine translation (NMT,~\citealp{luong2015effective,wu2016google,gehring2017convolutional,transformer}) where it benefits from a massive amount of training data. But obtaining such large amounts of parallel data is not-trivial in most machine translation scenarios. For example, there are many low-resource language pairs (e.g. English-to-Tamil), which lack adequate parallel data for training. 

Although many approaches about fully exploiting the parallel and monolingual data are proposed, e.g. back translation~\cite{sennrich-etal-2016-improving}, knowledge distillation~\cite{kim-rush-2016-sequence} and data diversification~\cite{nguyen2019data}, the prerequisite of these approaches is to build a well-performed baseline model based on the parallel data. However, \newcite{koehn2017six,lample2018phrase,sennrich2019revisiting} empirically reveal that NMT runs worse than their statistical or even unsupervised counterparts in low-resource conditions. Here naturally arise a question: \textit{Can we find a strategy to consistently improve NMT performance given the parallel data merely?}

We decide to find a solution from \textit{human learning behavior}.~\newcite{pavlenko2002bidirectional,dworin2003insights,chen2015bilingual} show that bidirectional language learning helps master bilingualism. In the context of machine translation, both the source$\rightarrow$target and target$\rightarrow$source language mappings may benefit bilingual modeling, which motivates many recent studies, e.g. dual learning~\cite{he2016dual} and symmetric training~\cite{cohn2016incorporating,Liang2007AgreementBasedL}. However, their approaches rely on external resources (e.g. word alignment or monolingual data) or complicated model modifications, which limit the applicability of the method to a broader range of languages and model structures. Accordingly, we turn to propose a simple data manipulation strategy and transfer the bidirectional relationship through \textit{bidirectional training} (\S\ref{subsec:bidirectional}). The core idea is using a bidirectional system as an initialization for a unidirectional system.
Specifically, to make the most of the parallel data, we first reconstruct the training samples from ``$\vec{\text{B}}$: source$\rightarrow$target'' to ``$\bid{\text{B}}$: source$+$target$\rightarrow$target$+$source'', where the training data was doubled. Then we update the model parameters with $\bid{\text{B}}$ in the early stage, and tune the model with normal ``$\vec{\text{B}}$ source$\rightarrow$target'' direction.

We validated our approach on several benchmarks across different language families and data sizes, including IWSLT21 En$\leftrightarrow$De, WMT16 En$\leftrightarrow$Ro, 
WMT19 En$\leftrightarrow$Gu,
IWSLT21 En$\leftrightarrow$Sw, WMT14 En$\leftrightarrow$De, WMT19 En$\leftrightarrow$De, WMT17 Zh$\leftrightarrow$En and WAT17 Ja$\leftrightarrow$En. Experimental results show that the proposed bidirectional training (BiT) consistently and significantly improves the translation performance over the strong Transformer~\cite{transformer}. 
Also, we show that BiT can complement existing data manipulation strategies, i.e. back translation, knowledge distillation and data diversification. Extensive analyses in \S\ref{subsec:analysis} confirm that the performance improvement indeed comes from the better cross-lingual modeling and our method works like a novel code-switching method.

\section{Bidirectional Training}
\subsection{Preliminary}
\label{subsec:standard}
Given a source sentence $\bf x$, an NMT model generates each target word ${\bf y}_t$ conditioned on previously generated ones ${\bf y}_{<t}$. Accordingly, the probability of generating $\bf y$ is computed as:
\begin{equation}
    p({\bf y}|{\bf x})
    =\prod_{t=1}^{T}p({\bf y}_t|{\bf x},{\bf y}_{<t}; \theta)
    \label{eq:standard}
\end{equation}
where $T$ is the length of the target sequence and the parameters $\theta$ are trained to maximize the likelihood of a set of training examples according to $\mathcal{L}(\theta) = \argmax_{\theta} \log p({\bf y}|{\bf x}; \theta)$. Typically, we choose Transformer~\cite{transformer} as its SOTA performance. The training examples can be formally defined as follows:
\begin{equation}
    \vec{\text{B}} = \{(\mathbf{x}_i, \mathbf{y}_i)\}^N_{i=1}
    \label{eq:B}
\end{equation}
where $N$ is the total number of sentence pairs in the training data. Note that in standard MT training, the $\bf{x}$ is fed into the encoder and $\textbf{y}_{<t}$ into the decoder to finish the conditional estimation for $\textbf{y}_t$, thus the utilization of $\vec{\text{B}}$ is directional, i.e. $\mathbf{x}_i$$\rightarrow$$\mathbf{y}_i$.

\subsection{Pretraining with Bidirectional Data}
\label{subsec:bidirectional}
\paragraph{Motivation}
The motivation is when human learn foreign languages with translation examples, e.g. $\mathbf{x}_i$ and $\mathbf{y}_i$. Both directions of this example, i.e. $\mathbf{x}_i$$\rightarrow$$\mathbf{y}_i$ and $\mathbf{y}_i$$\rightarrow$$\mathbf{x}_i$, may help human easily master the bilingual knowledge. 
Motivated by this, \newcite{levinboim-etal-2015-model,Liang2007AgreementBasedL} propose to modelling the invertibility between bilingual languages. \newcite{cohn2016incorporating} introduce extra bidirectional prior regularization to achieve symmetric training from the point view of training objective. \newcite{he2018layer,zheng2019mirror,ding2020acl} enhance the coordination of bidirectional corpus with model level modifications. Different from above methods, we model both directions of a given training example by a simple data manipulation strategy.

\paragraph{Our Approach}
Many studies have shown that pretraining could transfer the knowledge and data distribution, hence improving the generalization~\cite{hendrycks2019using,mathis2021pretraining}. Here we want to transfer the bidirectional knowledge among the corpus. Specifically, we propose to first pretrain MT models on bidirectional corpus, which can be defined as follows:
\begin{equation}
    \bid{\text{B}} = \{(\mathbf{x}_i, \mathbf{y}_i) \cup (\mathbf{y}_i, \mathbf{x}_i)\}^N_{i=1}
    \label{eq:bi.B}
\end{equation}
such that the $\theta$ in Equation~\ref{eq:standard} can be updated by both directions. Then the bidirectional pretraining objective can be formulated as:
\begin{align}
    \bid{\mathcal{L}}(\theta) = &\overbrace{\argmax_{\theta} \log p({\bf y}|{\bf x}; \theta)}^{\text{Forward}: \vec{\mathcal{L}_{\theta}}}\\
    &+ \underbrace{\argmax_{\theta} \log p({\bf x}|{\bf y}; \theta)}_{\text{Backward}: \cev{\mathcal{L}_{\theta}}}
    \label{eq:bi.B}
\end{align}
where the $\text{forward}~\vec{\mathcal{L}_{\theta}}$ and $\text{backward}~\cev{\mathcal{L}_{\theta}}$ are optimized iteratively. 

From data perspective, we achieve the bidirectional updating as follows: 1) swapping the source and target sentences of a parallel corpus, and 2) appending the swapped data to the original. Then the training data was doubled to make better and full use of the costly bilingual corpus.
The pretraining can acquire general knowledge from bidirectional data, which may help {\em better} and {\em faster} learning further tasks. 
Thus, we early stop BiT at 1/3 of the total training steps (we discuss its reasonability in \S\ref{subsec:setup}). In order to ensure the proper training direction, we further train the pretrained model on required direction $\vec{\text{B}}$ with the rest of 2/3 training steps.
Considering the effectiveness of pretraining~\cite{mathis2021pretraining} and clean finetuning~\cite{wu2019exploiting}, we introduce a combined pipeline: $\bid{\text{B}}\rightarrow\vec{\text{B}}$ as out best training strategy. There are many possible ways to implement the general idea of bidirectional pretraining. The aim of this paper is not to explore the whole space but simply to show that one fairly straightforward implementation works well and the idea is reasonable.

\begin{table*}[t]
    \centering
    \setlength{\tabcolsep}{3.2pt}
    \scalebox{0.98}{
    \begin{tabular}{lllllllllllc}
    \toprule
    \textbf{Data Source} &
    \multicolumn{2}{c}{\textbf{IWSLT14}} &
    \multicolumn{2}{c}{\textbf{WMT16}}  &
    \multicolumn{2}{c}{\textbf{IWSLT21}} & 
    \multicolumn{2}{c}{\textbf{WMT14}} & 
    \multicolumn{2}{c}{\textbf{WMT19}} &
    \multirow{2}{*}{$\Delta$}\\
    \textbf{Size} &
    \multicolumn{2}{c}{160K} &
    \multicolumn{2}{c}{0.6M} &
    \multicolumn{2}{c}{2.4M} &
    \multicolumn{2}{c}{4.5M} &
    \multicolumn{2}{c}{38M} &\\
    \cdashline{2-11}
    \textbf{Direction} & \bf En-De &\bf De-En & \bf En-Ro & \bf Ro-En  & \bf En-Sw & \bf Sw-En & \bf En-De &\bf De-En &\bf En-De &\bf De-En& \textbf{\textit{Ave.}}\\
    \midrule
    {\bf Transformer} & 29.2 & 35.1  & 33.9& 34.1  & 28.8 & 48.5 & 28.6 & 32.1 & 39.9 & 40.1 &--\\
    {\bf ~~~~+BiT} & 29.9$^\dagger$ & 36.3$^\ddagger$ & 35.2$^\ddagger$ & 35.9$^\ddagger$ & 29.9$^\ddagger$ & 49.9$^\ddagger$ & 29.7$^\ddagger$ & 32.9$^\dagger$ & 40.5$^\dagger$ & 41.6$^\ddagger$ & +1.1\\
    \bottomrule
    \end{tabular}}
    \caption{Comparison with previous AT work on several widely-used benchmarks, including IWSLT14 En$\leftrightarrow$De, WMT16 En$\leftrightarrow$Ro, IWSLT21 En$\leftrightarrow$Sw, WMT14 En$\leftrightarrow$De and WMT19 En$\leftrightarrow$De. ``$^{\ddagger/\dagger}$'' indicates significant difference ($p < 0.01/ 0.05$) from corresponding baselines, and this leaves as default symbol in Table~\ref{tab:distantlanguage}-\ref{tab:extreme-low}.}
    \label{tab:main-results}
\end{table*}
\section{Experiments}
\subsection{Setup}
\label{subsec:setup}
\paragraph{Data}
Main experiments in Table~\ref{tab:main-results} are conducted on five translation datasets: IWSLT21 English$\leftrightarrow$German~\cite{nguyen2019data}, WMT16 English$\leftrightarrow$Romania~\cite{gu2018non}, IWSLT21 English$\leftrightarrow$Swahili\footnote{\url{https://iwslt.org/2021/low-resource}}, WMT14 English$\leftrightarrow$German~\cite{transformer} and WMT19 English$\leftrightarrow$German\footnote{\url{http://www.statmt.org/wmt19/translation-task.html}}. The data sizes can be found in Table~\ref{tab:main-results}, ranging from 160K to 38M. 
Two distant language pairs in Table~\ref{tab:distantlanguage} are WMT17 Chinese$\leftrightarrow$English~\cite{hassan2018achieving} and WAT17 Japanese$\rightarrow$English~\cite{morishita2017ntt}, containing 20M and 2M training examples, respectively. The monolingual data used for back translation in Table~\ref{tab:complementary} is randomly sampled from publicly available News Crawl corpus\footnote{\url{http://data.statmt.org/news-crawl/}}.
We use same valid\& test sets with previous works for fair comparison except IWSLT21 English$\leftrightarrow$Swahili, where we follow \citet{Ding2021iwslt} to sample 5K/ 5K sentences from the training set as valid/ test sets.
We preprocess all data via BPE~\cite{Sennrich:BPE} with 32K merge operations. We use tokenized BLEU~\cite{papineni2002bleu} as the evaluation metric for all languages except English$\rightarrow$Chinese, where we use SacreBLEU\footnote{BLEU+case.mixed+lang.en-zh+numrefs.1+smooth.exp+test.wmt17+tok.zh+version.1.5.1}~\cite{post-2018-call}. The \textit{sign-test}~\cite{collins2005clause} is used for statistical significance test.

\paragraph{Model}
We validated our proposed BiT on Transformer~\cite{transformer}\footnote{\url{https://github.com/pytorch/fairseq}}. All language pairs are trained on Transformer-\textsc{Big} except IWSLT14 En$\leftrightarrow$De and WMT16 En$\leftrightarrow$Ro (trained on Transformer-\textsc{Base}) because of their extremely small data size. For fair comparison, we set beam size and length penalty as 5 and 1.0 for all language pairs.
It is worth noting that our data-level approach neither modifies model structure nor adds extra training loss, thus it's feasible to deploy on any frameworks, e.g. DynamicConv~\cite{wu2019pay} and non-autoregressive MT~\cite{gu2018non,Ding2020coling,ding2021iclr}, and training orders, e.g. curriculum learning~\cite{liu2020acl,zhou2021self,zhan2021meta,Ding2021acl2}. We will explore them in the future works. 

\paragraph{Training}
For Transformer-\textsc{Big} models, we empirically adopt large batch strategy~\cite{edunov2018understanding} (i.e. 458K tokens/batch) to optimize the performance. The learning rate warms up to $1\times10^{-7}$ for 10K steps, and then decays for 30K (data volumes range from 2M to 10M) / 50K (data volumes large than 10M) steps with the cosine schedule; For Transformer-\textsc{Base} models, we empirically adopt 65K tokens per batch for small data sizes, e.g. IWSLT14 En$\rightarrow$De and WMT16 En$\rightarrow$Ro. The learning rate warms up to $1\times10^{-7}$ for 4K steps, and then decays for 26K steps.
For regularization, we tune the dropout rate from [0.1, 0.2, 0.3] based on validation performance, and apply weight decay with
0.01 and label smoothing with $\epsilon$ = 0.1. We use Adam optimizer ~\citep{kingma2015adam} to train the models. We evaluate the performance on the averaged last 10 checkpoints to avoid stochasticity.

Someone may doubt that BiT heavily depends on how to properly set the early-stop steps. To dispel the doubt, we investigate whether our approach is robust to different early-stop steps. In preliminary experiments, we tried several simple fixed early-stop steps according to the size of training data (e.g. training 40K En-De and early stop at 10K/ 15K/ 20K, respectively). We found that both strategies achieve similar performances. Thus, we decide to choose a simple and effective method (i.e. 1/3 of the total training steps) for better reprehensibility.

\subsection{Results}
\paragraph{Results on Different Data Scales}
To confirm the effectiveness of our method across different data sizes,
we experimented on 10 language directions, including IWSLT14 En$\leftrightarrow$De, WMT16 En$\leftrightarrow$Ro, IWSLT21 En$\leftrightarrow$Sw, WMT14 En$\leftrightarrow$De and WMT19 En$\leftrightarrow$De. The smallest one merely contains 160K sentences, while the largest direction includes 38M sentence pairs. Table~\ref{tab:main-results} reports the results, we show that BiT achieves significant improvements over strong baseline Transformer in 7 out of 10 directions under the significance test $p<0.01$, and the rest of 3 directions also show promising performance under the significance test $p<0.05$, demonstrating the effectiveness and universality of our proposed bidirectional pretraining strategy.  
Notably, one advantage of BiT is it saves 1/3 of the training time for the reverse direction. For example, the pretrained BiT checkpoint for En$\rightarrow$De can be used to tune the reverse direction De$\rightarrow$En. This advantage shows BiT could be an efficient training strategy for multilinguality, e.g. multi-lingual pretraining~\cite{Liu2020mbart} and translation~\cite{hatoward}.

\paragraph{Results on Distant Language Pairs}
Inspired by \citet{Ding2021acl1}, to dispel the doubt that BiT could merely be applied on languages within the same language family, e.g. English and German, we report the results of BiT on Zh$\leftrightarrow$En and Ja$\rightarrow$En language pairs, which belong to different language families (i.e. Indo-European, Sino-Tibetan and Japonic). 

Table~\ref{tab:distantlanguage} lists the results, as seen, compared with baselines, our method significantly and incrementally improves the translation quality in all cases. In particular, BiT achieves averaged +0.9 BLEU improvement over the baselines, showing the effectiveness and universality of our method across language pairs.

\begin{table}[t]
    \centering
    \scalebox{1}{
    \begin{tabular}{lccc}
    \toprule
    \bf Data Source &   \multicolumn{2}{c}{\bf WMT17}   &   \multicolumn{1}{c}{\bf WAT17}  \\
    \bf Size    & \multicolumn{2}{c}{20M} &  \multicolumn{1}{c}{2M}\\
    \cdashline{2-3} \cdashline{4-4}
    \bf Direction  & \bf Zh-En &\bf En-Zh &\bf Ja-En\\
    \midrule
    {\bf Transformer} & 23.7 & 33.2 & 28.1 \\
    \bf {\bf ~~~~+BiT}& 24.9$^\ddagger$ & 33.9$^\dagger$ & 28.8$^\dagger$ \\
    \bottomrule 
    \end{tabular}}
    \caption{Performance on distant language pairs, including WMT17 Zh$\leftrightarrow$En and WAT17 Ja$\rightarrow$En. To perform BiT on languages in different alphabets, we share the sub-words dictionaries between languages.}
    \label{tab:distantlanguage}
\end{table}

\begin{table}[t]
    \centering
    \scalebox{0.95}{
    \begin{tabular}{lc}
    \toprule
    \bf Model & \bf BLEU \\
    \midrule
    \textbf{Transformer-\textsc{Big}}/+BiT & 28.6/ 29.7$^\ddagger$\\
    \hdashline
    \textbf{~~+BT}\cite{caswell-etal-2019-tagged}/+BiT & 30.5/ 31.2$^\dagger$\\
    \hdashline
    \textbf{~~+KD}\cite{kim-rush-2016-sequence}/+BiT & 29.3/ 30.1$^\dagger$\\
    \hdashline
    \textbf{~~+DD}\cite{nguyen2019data}/+BiT & 30.1/ 30.7$^\dagger$\\
    \bottomrule
    \end{tabular}}
\caption{Complementary to other works. ``/+BiT'' means combining BiT with corresponding works, and BLEU scores of BiT followed their counterparts with ``/''. Experiments are conducted on WMT14 En-De.}
\label{tab:complementary}
\end{table}

\paragraph{Complementary to Related Work}
Recent studies start to combine pretraining and traditional data manipulation approaches for better model performance~\citep{lample2019cross,Liu2020mbart,liu2021emnlp}. To show the complementary between our proposed pretraining method BiT and related data manipulation works,
we list three representative data manipulation approaches for NMT: a) {Tagged Back Translation} (\textbf{BT},~\citealt{caswell-etal-2019-tagged}) combines the synthetic data generated with target-side \textit{monolingual data} and parallel data; b) {Knowledge Distillation} (\textbf{KD},~\citealt{kim-rush-2016-sequence}) trains the model with sequence-level distilled \textit{parallel data}; c) {data diversification} (\textbf{DD},~\citealt{nguyen2019data}) 
diversifies the data by applying KD and BT on \textit{parallel data}. 
As seen in Table~\ref{tab:complementary}, BiT can be applied on existing data manipulation approaches and yield further significant improvements. 

\subsection{Analysis}
\label{subsec:analysis}
We conducted analyses to better understand BiT. Unless otherwise stated, all results are reported on the WMT14 En-De.
\begin{table}[tb]
    \centering
    \scalebox{1}{
    \begin{tabular}{ll}
        \toprule
         \textbf{Model} & \bf BLEU  \\
        \midrule
         \bf Transformer-\textsc{Big}&28.6\\
         {\bf ~~~+mRASP}~\cite{lin2020pre}&29.3$^\dagger$\\
         {\bf ~~~+CSP}~\cite{yang2020csp}&29.4$^\dagger$\\
         {\bf ~~~+BiT} (Ours)&29.7$^\ddagger$\\
        \bottomrule
    \end{tabular}}
    \caption{Comparison with previous code-switch approaches on bilingual data, where we follow the best settings of ``+mRASP'' and ``+CSP'' as default without extra parameter tuning. For fair comparison, the pretrain/ finetune steps are identical with ours.}
    \label{tab:codeswitch}
\end{table}
\paragraph{BiT works as a simple bilingual code-switcher}
\newcite{lin2020pre,yang2020csp} employ the third-party tool to obtain the alignment information to perform code-switching pertraining, where partial of the source tokens is replaced with the aligned target ones. But training such alignment model is time-consuming and the alignment errors may be propagated. 
Actually, BiT can be viewed as a novel yet simple bilingual code-switcher, where the switch span is the whole sentence and both the source- and target-side sentences are replaced with the probability $0.5$. 
\begin{CJK}{UTF8}{gbsn}
Take a sentence pair \{``Bush held a talk with Sharon''$\rightarrow$``布什 与 沙龙 举行 了 会谈''\} in English$\rightarrow$Chinese dataset as an example, during pretraining phase, the reconstructed corpus contains \{``Bush held a talk with Sharon'' $\rightarrow$ ``布什 与 沙龙 举行 了 会谈''\} and its reversed version {``布什 与 沙龙 举行 了 会谈'' $\rightarrow$ ``Bush held a talk with Sharon''}, simultaneously. For the English$\rightarrow$Chinese direction, the reversed sentence pair exactly belongs to the sentence-level switch with a probability of 0.5. \end{CJK}
For fair comparison, we implement \citealp{lin2020pre,yang2020csp}'s approaches in \textit{bilingual data scenario}. Table~\ref{tab:codeswitch} show the superiority of BiT, indicating BiT is a good alternative to code-switch in bilingual scenario.

\paragraph{BiT improves alignment quality}
Our proposed BiT intuitively encourages self-attention to learn bilingual agreement, thus has the potential to induce better attention matrices. 
We explore this hypothesis on the widely-used Gold Alignment dataset\footnote{\url{http://www-i6.informatik.rwth-aachen.de/goldAlignment}, the original dataset is German-English, we reverse it to English-German.} and follow~\citet{tang2019understanding} to perform the alignment. The only difference being that we average the attention matrices across all heads from the penultimate layer~\cite{garg2019jointly}. The alignment error rate (AER,~\citealt{och2003systematic}), precision (P) and recall (R) are evaluation metrics. Table~\ref{tab:align} summarizes that BiT allows Transformer to learn better attention matrices, thereby improving alignment performance (24.3 vs. 27.1).
\begin{table}[tb]
    \centering
    \scalebox{0.97}{
    \begin{tabular}{lccc}
        \toprule
         \textbf{Model}&{\textbf{AER}}&\textbf{P}&\textbf{R}  \\
        \midrule
         \bf Transformer-\textsc{Big}&27.1\%&71.2\%&74.7\%\\
         {\bf ~~~~+BiT}&24.3\%&74.6\%&76.9\%\\
        \bottomrule
    \end{tabular}}
    \caption{The AER scores of alignments on En-De.}
    \label{tab:align}
\end{table}

\begin{table}[tb]
    \centering
    \scalebox{1}{
    \begin{tabular}{lccc}
        \toprule
         \textbf{Model}&{\textbf{En$\rightarrow$Gu}}&\textbf{Gu$\rightarrow$En}&\textbf{Ave.$\Delta$}  \\
        \midrule
         \bf Base & 3.0 & 8.2 & -\\
         \rowcolor{liam}{\bf ~~~~+BT}& 2.6 & 10.1 & - \\
         \hdashline
         \bf Base+BiT & 4.2$^\ddagger$ & 9.0$^\ddagger$ & +1.0\\
         \rowcolor{liam}{\bf ~~~~+BT}& 5.8$^\ddagger$ & 12.4$^\ddagger$ & +2.8\\
        \bottomrule
    \end{tabular}}
    \caption{Results for En$\leftrightarrow$Gu on WMT2019 test sets. ``Ave. $\Delta$'' shows the averaged improvements of ``Base+BiT'' v.s. ``Base'' and their corresponding ``+BT'' comparisons.}
    \label{tab:extreme-low}
\end{table}

\paragraph{BiT works for extremely low-resource settings} 
Researches may doubt BiT may fail on extremely low-resource settings where \textit{back-translation even does not work}. To dispel this concern, we conduct experiments on WMT19 English$\leftrightarrow$Gujurati\footnote{\url{http://www.statmt.org/wmt19/translation-task.html}} in Table~\ref{tab:extreme-low}. Specifically, we follow \citet{li2019niutrans} to collect and preprocess the parallel data to build the base model ``Base'' and our ``Base+BiT'' model. For a fair comparison, we sample the monolingual English and Gujurati sentences to ensure Parallel: Monolingual = 1:1 to generate the synthetic data. As seen, when directly applying back-translation (BT) on the En$\leftrightarrow$Gu Base model, there indeed shows a slight performance drop (-0.4 BLEU). However, our ``BiT'' significantly improves the initial Base model by averaged +1.0 BLEU, and making the BiT-equipped BT more effective compared to vanilla BT (+2.8 BLEU). 
These findings on extremely low-resource settings demonstrate that 1) our BiT consistently works well; and 2) BiT provides a better initial model, thus rejuvenating the effects of back-translation.

\section{Conclusion and Future Works}
In this study, we propose a pretraining strategy for NMT with parallel data merely. Experiments show that our approach significantly improves translation performance, and can complement existing data manipulation strategies. Extensive analyses reveal that our method can be viewed as a simple yet better bilingual code-switching approach, and improves bilingual alignment quality. 

Encouragingly, with BiT, our system~\cite{Ding2021iwslt} got the first place in terms of BLEU scores in IWSLT2021\footnote{\url{https://iwslt.org/2021/}} low-resource track. 
It will be interesting to
integrate BiT into our previous systems~\cite{ding2019wmt,Wang2020TencentAL} and validate its effectiveness on industrial level competitions, e.g. WMT\footnote{\url{http://www.statmt.org/wmt21/}}. It is also worthwhile to explore the effectiveness of our proposed bidirectional pretraining strategy on multilingual NMT task~\cite{hatoward}.
\section*{Acknowledgments}
We would thank the anonymous reviewers and the area chair for their considerate proofreading and valuable comments.

\bibliography{custom}
\bibliographystyle{acl_natbib}

\appendix
\end{document}